OpenStudy (anonymous):

HELP WITH TWO QUESTIONS FOR FAN + MEDAL!!! 1. Estimate the solution to the equation by first rounding the constants and coefficients to the nearest integer. 6.15(x – 8.86) = 21.83 x ≈ ____ 2. Michaela purchases a box of cereal for $3.95, chicken for $5.99, orange juice for $2.10, and a loaf of bread for $1.05. She has $15 with her. About how much change should she expect to receive? Round each amount to the nearest dollar to find your estimate. $____

OpenStudy (anonymous):

i can help u

OpenStudy (anonymous):

1:question rounding gives: 6(x-9) = 22 dividing by 6 gives and adding 9 gives: x = 12.66

OpenStudy (anonymous):

srry i'm having technical problems and thanks :) do u know the second one?

OpenStudy (anonymous):

2 question 15 - 3.95 - 5.99 - 2.10 - 1.05 = change change = $1.96 ≈ $2

OpenStudy (anonymous):

i think

OpenStudy (anonymous):

i try to help u dont be mad please

OpenStudy (anonymous):

#2 Add the prices together 3.95+5.99+2.10+1.05=13.09 Round 13.09 to the nearest integer. so u would get 13

OpenStudy (anonymous):

13-15=2 in change

OpenStudy (anonymous):

that help

OpenStudy (anonymous):

ohhh well i got it wrong but i did the first one

OpenStudy (anonymous):

how do i know who's right?

OpenStudy (anonymous):

ohhh i got it write yay

OpenStudy (anonymous):

I'll fan both of u but i don't know who to medal...

OpenStudy (anonymous):

dont worry medal him

OpenStudy (anonymous):

im nice guy

OpenStudy (anonymous):

u there